Required information [The following information applies to the questions displayed below.) Data for Hermann Corporation are sReq 1A Req 1B Should the advertising budget be increased? Yes ONORequired information [The following information applies to the questions displayed below.) Data for Hermann Corporation are sReq 2A Req 2B Should the higher-quality components be used? Yes O No
Answer #1

1a) Increase profit (loss) = Incremental revenue-Incremental cost

= (21000*40%)-9200

Increase profit (loss) = -800

1b) No

2a) Present net income = (2900*54)-87000 = 69600

proposed net income = (2900*1.20*49)-87000 = 83520

Net operating income increase by (83520-69600) = 13920

2b) Yes